OpenStudy (anonymous):

Which operation would isolate the variable in this equation ? 3y-12=14

OpenStudy (kainui):

Add twelve to both sides, and then divide both sides by 3. Try it out.

OpenStudy (uri):

Add 12 both sides 3y-12=14 +12 12 ========== 3y=26 divide 26 wih 3 y=26/3 y=8.66